If (2, 3, 5) is one end of a diameter of the sphere x2 + y2 + z2 – 6x – 12y – 2z + 20 = 0, then the cooordinates of the other end of the diameter ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Equation of sphere is x2+y2+z2-6x-12y-2z+20=0
Whose coordinates of centre are (3,6,1)
Let the coordinates of other end of diameter are (a,b,c)
then (a+2)/2=3; (b+3)/2=6 ; (c+5)/2=1
we get a=4,b=9,c=-3
Thus coordinates of other point are (4,9,-3)
Thanks and Regards
Shaik Aasif
askIITians faculty